simplify:
(x^4)^-3 * 2x^4
1
Expert's answer
2012-09-11T11:12:25-0400
(x^4)^-3 * 2x^4= 2x^4/{(x^4)^3}=2/{(x^4)}^2=2/(x^8)=2x^{-8}